OpenStudy (anonymous):

Half angle identity for cos 13pi / 12

OpenStudy (anonymous):

in degree or radian

OpenStudy (mathmale):

"Half angle identity for cos 13pi / 12 " comes across as vague? I'm not really sure what yu want to calculate/determine/find. Would you mind going back to the original source of this problem statement and double check to ensure that you've included everything?

OpenStudy (anonymous):

cos (13 Π /12 ) = cos [(12 Π + Π)/12 ] = cos(Π + Π/12) = - cos(Π/12) = - cos [Π/3 - Π/4 ] = - [ cos (Π/3) cos(Π/4) + sin(Π/3)sin(Π/4) = - [ (1/2)(1 /√2) + (√3/2)(1/√2) ] = - [ 1 + √3 ] / 2√2 = - (1/4)(√2 + √6)

OpenStudy (anonymous):

= -(√6 + √2)\8 is the half angel

OpenStudy (mathmale):

I suspect that you were asked to use the "half angle formula for the cosine" in arriving at a solution to this problem (finding the cos of the given angle). Please see the attached illustration, which comes from http://www.purplemath.com/modules/idents.htm.

OpenStudy (mathmale):

The cosines of pi/2, pi/3, pi/6, and so on, are easily found, Not so easily found is the cosine of pi/12. this is the reason for using the "half angle formula" for the cosine. Try this, please.